Show that the fugacity of a van der


Waals gas is given by the expression.



In f = -In (v-b) - 2a/RTV + b/v-b

1
Expert's answer
2021-12-13T10:34:52-0500

"\ln f = -\ln (v-b) - \frac{2a}{RTV }+\frac b{v-b}."